OpenStudy (anonymous):

if r1 and r2 are the roots of the quadratic equation ax+bx+c=0 show that r1+r2=-b/and r1r2=c/a

OpenStudy (mertsj):

r1+r2 does not equal -b so it is futile to attempt to show that it does.

OpenStudy (anonymous):

typo show that r1+r2=-b/a and r1 r2=c/a

OpenStudy (mertsj):

Are we to assume that you also meant ax^2+bx+c=0?

OpenStudy (anonymous):

sorry let me retype the question

OpenStudy (anonymous):

It has been one of those days

OpenStudy (mertsj):

We all have them.

OpenStudy (anonymous):

if r1 and r2 are the root of the quadratic equation ax^2+bx=c=0 shows thar r1+r2=-b/a and r1r2=c/a

OpenStudy (campbell_st):

make the quadratic monic by dividing by a x^2 + b/ax + c/a = 0 let \[r _{1} r _{2} be the roots of (x - r _{1})(x- r _{2})\] expanding \[x^2 -(r _{1}+r _{2})x + r _{1} \times r _{2}\] equating coefficients of the quadratics then \[a = 1, \] \[b/a = -(r _{1} + r _{2})\] a little number theory allows \[-b/a = r _{1} + r _{2}\] the Sum \[c/a = r _{1} \times r _{2}\] the product

OpenStudy (anonymous):

that is like major confusing, But the more I do the more i hope to get it Thanks

OpenStudy (campbell_st):

just remember... sum = -b/a and product is c/a.... this is a theoretical question to see if you can do algebra....

OpenStudy (anonymous):

that would make sense